# A Dual-way Enhanced Framework from Text Matching Point of View for Multimodal Entity Linking

### v1

@@ Introduction
Dozens of MEL studies have been conducted in the past few years and achieved promising performances, where delicate interaction and fusion mechanisms are designed for encoding the image and text features (Wang, Wu, and Chen 2022; Zhou et al. 2021; Chen et al. 2022b). Nonetheless, existing approaches still struggle to fully exploit the potential of the feature sources from two distinct information perspectives, which may hinder further task development.
@@
Ambiguous Entity Representation: Wang et al. (2022) and Zhou et al. (2021) concatenate the main properties of an entity into a piece of text, which is regarded as a representation of the entity in KG. Sun et al.(2022b) manually collect short text to represent entity, e.g., “Bahador Abb: Iranian footballer”. However, the set of properties or short text are insufficient to represent entities. Different people may have the same properties. For example, the entity Donald Trump and Mike Pompeo both have the properties of “Sex: male. Religion: Presbyterianism. Occupation: businessperson, politician. Languages: English. work location: Washington, D.C.” With so similar entity representation, it is a great challenge to link entity even with an ideal multimodal method. Thus, more distinctive and representative entity representation is required for further disambiguation. Fortunately, Wikipedia description offers a promising solution, which has been shown to enrich the semantics of the raw data, especially in low-resources NLP tasks (Chen et al. 2022a).
@@
Extensive experiments show that DWE pushes the state-of-the-art (SOTA) by a large margin, indicating the efficiency of DWE. To summarize, our contributions are as follows:
@@
Knowledge base Enhanced Entity Representation. Previous works (Wang, Wu, and Chen 2022; Zhou et al. 2021) tend to concatenate the main properties of an entity into a piece of text and regard the properties as textual entity representation. However, a set of properties is not clear enough to distinguish the linking entity from candidate entities. Different entity may be the same in some fields. As shown in the entity representation in Figure 2, the properties of Mike Pompeo is similar with Donald Trump in previous representation. Thus it is challenging to distinguish the two people with nearly the same properties.
@@
(1) BLINK (Wu et al. 2019) is an BERT-based EL model with a two-stage zero-shot linking. (2) BERT (Devlin et al. 2019) stacks several layers of transformer to encode each token in text. (3) ARNN (Eshel et al. 2017) utilizes Attention-RNN to predict associations with candidate entity textual features. (4) DZMNED (Moon, Neves, and Carvalho 2018) takes concatenated multimodal attention mechanism to fuse visual, textual, and character features of mentions. (5) JMEL (Adjali et al. 2020b) utilizes fully connected layers to project the visual and textual features into an implicit joint space. (6) MEL-HI (Zhang, Li, and Yang 2021) uses multiple attention to get richer information and refine the negative impact of noisy image. (7) HieCoAtt (Lu et al. 2016b) is a multimodal fusion mechanism, using alternating co-attention and three textual levels (tokens, phrases, and sentences) to calculate co-attention maps. (8) GHMFC (Wang, Wu, and Chen 2022) is a advanced baseline proposed by Wang et.al (Wang, Wu, and Chen 2022). It takes the gated multimodal fusion and novel attention mechanism to link multimodal entities. (9) MMEL (Yang et al. 2023) is a joint feature extraction module to learn the representations of context and entity candidates, from both the visual and textual perspectives. (10) CLIP-text (Radford et al. 2021) only use textual information such as text and mention in the dataset. It mainly focuses on the ability to understand textual relationship among text, mention, and entity representation. (11) CLIP (Radford et al. 2021) take both textual and visual features into consideration. The model concatenates multimodal features and calculates the similarity between features and ground-truth.
@@
$m_{t}$
$m_{v}$
$\mathcal{L}_{c}$
$m_{s}$
$f$
Top-1
Top-5
Top-10
Top-20
0
72.5
97.3
98.8
99.6
1
70.9
97.5
98.8
99.5
2
69.8
97.3
98.4
99.4
3
67.3
93.9
98.4
99.4
4
63.5
92.7
96.0
98.3
Table 4: Ablation study on using refined visual feature $v$ or primitive visual feature which is extracted directly from image over two public datasets: Richpedia (rich) and WikiMEL (wiki)
@@
Enhanced Entity Representation. Previous works (Zhou et al. 2021; Wang, Wu, and Chen 2022) consider the properties as entity representation, which is ambiguous and limited. we take textual Wikipedia description as enhanced representation. Results show that our method is not only useful for previous entity representation but also the enhanced representation. Besides, on CLIP and our method, the improvement over different representations shows that the enhanced representation is helpful for disambiguation. In detail, on Richpedia, by using enhanced representation, CLIP obtain the 3.5% improvement and our method obtains 4.9% improvement. The boost proves that the enhanced entity representation makes the entity more distinctive.
@@
Table 6: Multimodal entity linking cases. Entity is depicted by textual Wikipedia description. Top-5 entities are selected from $\lambda$ candidate entities. We take the bold blue text to show the correct entity. Case 1 and 2 are multiple objects samples for different mentions with the same image and text.
Case
1 2 3 4
Image
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Text
Khanna and Manek at the music launch of Ra.One along with Shahrukh Khan, Pooja Gaur and Sanaya Irani
Khanna and Manek at the music launch of Ra.One along with Shahrukh Khan, Pooja Gaur and Sanaya Irani
Miller and Saxby Chambliss in 2004
Messi has scored over 700 senior career goals for club and country
Mention
Manek Khan Miller Messi
Entity
Giaa Manek is an Indian television actress. Before venturing into the television industry, she played a minor role in the 2010 Hindi comedy film Na Ghar Ke Na Ghaat Ke..
Shah Rukh Khan (born 2 November 1965), also known by the initialism SRK, is an Indian actor, film producer, and television personality who works in Hindi films..
Zell Bryan Miller (February 24, 1932 -2013 March 23, 2018) was an American author and politician from the state of Georgia..
Lionel Messi (born 24 June 1987), also known as Leo Messi, is an Argentine professional footballer who plays as a forward for Ligue 1 club Paris..
Top-5
candidate
Giaa Manek 0.91
Gia Kancheli 0.34
Mihir Bipin Manek 0.29
Brady Manek 0.28
Rama 0.04
Rajesh Khanna 0.76
Shah Rukh Khan 0.72
Zareen Khan 0.43
Gauhar Khan 0.33
Aamir Khan 0.19
Sienna Miller 0.79
Marisa Miller 0.61
Kristine Miller 0.56
Ezra Miller 0.31
Dean Miller 0.30
Lionel Messi 0.85
Antonello da Messina 0.76
Charles Messier 0.56
Massiv 0.29
Ramesses II 0.06
Prediction
@@
Attribute enhancement The visual attributes extracted from the image depict the entity’s outline and some prominent features. The extracted visual attributes from the image include facial features and adjective-noun pair (ANPs), which play different roles in the model. (1) Facial features: When the entity in the image is a human subject, the extracted facial features serve as a supplement and correction to the feature extractor, such as in cases where the feature extractor fails to recognize a white male in the image. As shown in Table 3, by removing the facial features (method 0$\rightarrow$1), the model’s performance decreased by 1.6%, demonstrating the effectiveness of facial features. (2) ANPs: When the image depicts non-human entities like places, ANPs are coarse-grained concepts that describe scenes, such as “nice clouds,” “white dress,” and “clear sky.” However, due to the presence of noise in ANPs, we use a gated fusion to integrate them. By removing (method 1$\rightarrow$2) attribute-enhanced mention $m_{s}$ (ANPs), the accuracy decreased by 1.1%.
@@
We first analyze the correct predictions. The correctness of case 1 and 4 shows the case with multiple visual objects on different datasets. As for the correct case 3, the mention is Miller with two old men in image. The result shows the effectiveness of the attributes. The model could mine the relationship between two old men, and entity representation of Sienna Miller , while the others are female. On the other hand, wrong predictions can still show the effectiveness of our model. As shown in case 2, The visual and textual information includes many objects such as Khan and Khanna. It is difficult to distinguish these two mentions by the same multimodal information. However, it could be seen that the similarity of Rajesh Khanna and Shah Rukh Khan is much larger than the rest of candidate entity, which still shows the learning ability of DWE across modalities.
@@ Conclusion
In this paper, we explore Multimodal Entity Linking in a neural matching formulation and design a dual-way enhanced framework: query is enhanced by refined multimodal information and the semantics of entity representation is enriched by Wikipedia. DWE utilizes a pretrained visual encoder and vision tools to acquire the visual representation and visual attributes, which are subsequently refined into learnable visual characteristics. It regards learnable visual characteristics and textual content as refined information, each contributing to the enhancement of the mention via three enhanced units. During the process, we apply the crossmodal alignment to bridge the semantic gap among visual and textual features to gain the multi-view alignment representation. Extensive experiments show that DWE outperforms the state-of-the-art methods on three public datasets: Richpedia, WikiMEL and Wikidiverse. Furthermore, we introduce an enhanced entity representation, i.e., textual Wikipedia description. In this way, the entity representation is more representative and no longer restricts the learning of multimodal information.
Acknowledgement

### v2

@@ Introduction
Dozens of MEL studies have been conducted in the past few years and achieved promising performances, where delicate interaction and fusion mechanisms are designed for encoding the image and text features (Wang 2022; Zhou et al. 2021; Chen et al. 2022b). Nonetheless, existing approaches still struggle to fully exploit the potential of the feature sources from two distinct information perspectives, which may hinder further task development.
@@
Ambiguous Entity Representation: Wang et al. (2022) and Zhou et al. (2021) concatenate the main properties of an entity into a piece of text, which is regarded as a representation of the entity in KG. Sun et al.(2022b) manually collect short text to represent entity, e.g., “Bahador Abb: Iranian footballer”. However, the set of properties or short text are insufficient to represent entities. Different people may have the same properties. For example, the entity Donald Trump and Mike Pompeo both have the properties of “Sex: male. Religion: Presbyterianism. Occupation: businessperson, politician. Languages: English. work location: Washington, D.C.” With so similar entity representation, it is a great challenge to link entity even with an ideal multimodal method. Thus, more distinctive and representative entity representation is required for further disambiguation. Fortunately, Wikipedia description offers a promising solution, which has been shown to enrich the semantics of raw data, especially in low-resources NLP tasks (Chen et al. 2022a).
@@
Extensive experiments show that DWE pushes the state-of-the-art (SOTA) by a large margin, indicating the efficiency of DWE. Our contributions are as follows:
@@
Knowledge base Enhanced Entity Representation. Previous works (Wang 2022; Zhou et al. 2021) tend to concatenate the main properties of an entity into a piece of text and regard the properties as textual entity representation. However, a set of properties is not clear enough to distinguish the linking entity from candidate entities. Different entity may be the same in some fields. As shown in the entity representation in Figure 2, the properties of Mike Pompeo is similar with Donald Trump in previous representation. Thus it is challenging to distinguish the two people with nearly the same properties.
@@
(1) BLINK (Wu et al. 2019) is an BERT-based EL model with a two-stage zero-shot linking. (2) BERT (Devlin et al. 2019) stacks several layers of transformer to encode each token in text. (3) ARNN (Eshel et al. 2017) utilizes Attention-RNN to predict associations with candidate entity textual features. (4) DZMNED (Moon, Neves, and Carvalho 2018) takes concatenated multimodal attention mechanism to fuse visual, textual, and character features of mentions. (5) JMEL (Adjali et al. 2020b) utilizes fully connected layers to project the visual and textual features into an implicit joint space. (6) MEL-HI (Zhang, Li, and Yang 2021) uses multiple attention to get richer information and refine the negative impact of noisy image. (7) HieCoAtt (Lu et al. 2016b) is a multimodal fusion mechanism, using alternating co-attention and three textual levels (tokens, phrases, and sentences) to calculate co-attention maps. (8) GHMFC (Wang 2022) is a advanced baseline proposed by Wang et.al (Wang 2022). It takes the gated multimodal fusion and novel attention mechanism to link multimodal entities. (9) MMEL (Yang et al. 2023) is a joint feature extraction module to learn the representations of context and entity candidates, from both the visual and textual perspectives. (10) CLIP-text (Radford et al. 2021) only use textual information such as text and mention in the dataset. It mainly focuses on the ability to understand textual relationship among text, mention, and entity representation. (11) CLIP (Radford et al. 2021) take both textual and visual features into consideration. The model concatenates multimodal features and calculates the similarity between features and ground-truth.
$m_{t}$ $m_{v}$ $\mathcal{L}_{c}$ $m_{s}$ $f$ Top-1 Top-5 Top-10 Top-20
0 72.5 97.3 98.8 99.6
1 70.9 97.5 98.8 99.5
2 69.8 97.3 98.4 99.4
3 67.3 93.9 98.4 99.4
4 63.5 92.7 96.0 98.3
@@
Table 4: Ablation study on using refined visual feature $v$ or primitive visual feature which is extracted directly from image over public datasets: Richpedia (rich), WikiMEL (wiki)
@@
Enhanced Entity Representation. Previous works (Zhou et al. 2021; Wang 2022) consider the properties as entity representation, which is ambiguous and limited. we take textual Wikipedia description as enhanced representation. Results show that our method is not only useful for previous entity representation but also the enhanced representation. Besides, on CLIP and our method, the improvement over different representations shows that the enhanced representation is helpful for disambiguation. In detail, on Richpedia, by using enhanced representation, CLIP obtain the 3.5% improvement and our method obtains 4.9% improvement. The boost proves that the enhanced entity representation makes the entity more distinctive.
@@
Table 5: Top-1 accuracy for the model from 16 candidates.
Case 1 2 3 4
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Text Khanna and Manek at the music launch of Ra.One Khanna and Manek at the music launch of Ra.One Miller and Saxby Chambliss in 2004 Messi has scored over 700 senior career goals
Mention Manek Khan Miller Messi
Entity Giaa Manek is an Indian television actress, who played a role in film Na Ghar Ke.. Shah Rukh Khan also known by the initialism SRK, is an Indian actor, film producer.. Zell Bryan Miller (1932-2013) was an American author and politician.. Lionel Messi, also known as Leo Messi, is an Argentine professional footballer..
Top-5 candidate Giaa Manek 0.91 Gia Kancheli 0.34 Mihir Bipin Manek 0.29 Brady Manek 0.28 Rama 0.04 Rajesh Khanna 0.76 Shah Rukh Khan 0.72 Zareen Khan 0.43 Gauhar Khan 0.33 Aamir Khan 0.19 Sienna Miller 0.79 Marisa Miller 0.61 Kristine Miller 0.56 Ezra Miller 0.31 Dean Miller 0.30 Lionel Messi 0.85 Antonello da Messina 0.76 Charles Messier 0.56 Massiv 0.29 Ramesses II 0.06
Predict
Table 6: Multimodal entity linking cases. Entity is depicted by textual Wikipedia description. Top-5 entities are selected from $\lambda$ candidate entities. We take the bold text to show the correct entity in Top-5 candidate. Case 1 and 2 are multiple object samples for different mentions with the same image and text.
@@
Attribute enhancement The visual attributes extracted from the image depict the entity’s outline and some prominent features. The extracted visual attributes from the image include facial features and adjective-noun pair (ANPs), which play different roles in the model. (1) Facial features: When the entity in the image is a human subject, the extracted facial features serve as a supplement and correction to the feature extractor, such as in cases where the feature extractor fails to recognize a white male in the image. As shown in Table 3, by removing the facial features (method 0$\rightarrow$1), the model’s performance decreased by 1.6%, demonstrating the effectiveness of facial features. (2) ANPs: When the image depicts non-human entities like places, ANPs are coarse-grained concepts that describe scenes, such as “nice clouds,” and “white dress.” Due to the presence of noise in ANPs, we use a gated fusion to integrate them. By removing (method 1$\rightarrow$2) attribute-enhanced mention $m_{s}$ (ANPs), the accuracy decreased by 1.1%.
@@
We first analyze the correct predictions. The correctness of case 1 and 4 shows the case with multiple visual objects on different datasets. As for the correct case 3, the mention is Miller with two old men in image. The result shows the effectiveness of the attributes. The model could mine the relationship between two old men, and entity representation of Sienna Miller , while the others are female. On the other hand, wrong predictions can still show the effectiveness of our model. In case 2, the visual and textual information includes many objects such as Khan and Khanna. It is difficult to distinguish these two mentions by the same multimodal information. However, it could be seen that the similarity of Rajesh Khanna and Shah Rukh Khan is much larger than the rest of candidate entity, which still shows learning ability.
@@ Conclusion
In this paper, we explore Multimodal Entity Linking in a neural matching formulation and design a dual-way enhanced framework: query is enhanced by refined multimodal information and the semantics of entity representation is enriched by Wikipedia. DWE utilizes a pretrained visual encoder and vision tools to acquire the visual representation and visual attributes, which are subsequently refined into learnable visual characteristics. It regards learnable visual characteristics and textual content as refined information, each contributing to the enhancement of the mention via three enhanced units. During the process, we apply the crossmodal alignment to bridge the semantic gap among visual and textual features to gain the multi-view alignment representation. Extensive experiments show that DWE outperforms the state-of-the-art methods on three public datasets. Furthermore, we introduce an enhanced entity representation, i.e., textual Wikipedia description. The enhanced entity representation is more representative and no longer restricts the learning of multimodal information.
Acknowledgements